OpenStudy (anonymous):

please any body can solve this.. lim x->0 cosecX-cotX/x

terenzreignz (terenzreignz):

\[\lim_{x \rightarrow 0}\frac{\csc x - \cot x}{x}\] Familiar with L'Hopital's Rule?

OpenStudy (anonymous):

whats is l hospital rule?

terenzreignz (terenzreignz):

on second thought, I may have been hasty... forget about that :) Just focus on csc x - cot x Notice anything?

OpenStudy (anonymous):

cant apply this..

terenzreignz (terenzreignz):

Yeah, I realised (a bit too late :) ) But focus on csc x - cot x

OpenStudy (anonymous):

whats i got?

terenzreignz (terenzreignz):

\[\csc x - \cot x = \frac{1}{\sin x} - \frac{\cos x}{\sin x}=\frac{1-\cos x}{\sin x}\] Look familiar?

OpenStudy (anonymous):

please tell me full steps how to solve this. i dont know L hostipal rule and i cant apply that

terenzreignz (terenzreignz):

Okay, we might not be able to use L'Hopital, but maybe we can be sneaky with it :) First things first, though, to make it simple, carrying on from that csc x - cot x bit: \[\frac{1-\cos x}{\sin x}=\sqrt{\frac{(1-\cos x)^{2}}{\sin^{2}x}}=\sqrt{\frac{(1-\cos x)(1-\cos x)}{1-\cos^{2}x}}\] \[=\sqrt{\frac{(1-\cos x)(1-\cos x)}{(1-\cos x)(1+\cos x)}}\] cancelling out 1-cos x we get\[=\sqrt{\frac{1-\cos x}{1+\cos x}}\] Does it look familiar now?

OpenStudy (anonymous):

ok than?

terenzreignz (terenzreignz):

\[\sqrt{\frac{1-\cos x}{1+\cos x}}=\tan \left( \frac{x}{2} \right)\] Can you do it from here?

OpenStudy (anonymous):

tell me please how can i solve that question..

terenzreignz (terenzreignz):

What have we done so far? We've shown that \[\csc x - \cot x = \tan \left( \frac{x}{2} \right)\] right?

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

For simplicity, we can let \[f(x)=\tan \frac{x}{2}\] so we can write \[\lim_{x \rightarrow 0}\frac{\csc x - \cot x}{x}=\lim_{x \rightarrow 0}\frac{f(x)}{x}\] Catch me so far?

OpenStudy (anonymous):

look i got \[\lim_{x \rightarrow o}{\frac{ 1 }{ \sin }-\frac{ cosx }{ sinx }}\]

OpenStudy (anonymous):

ok than?

terenzreignz (terenzreignz):

And you'd end up with an indeterminate form... which would require the use of L'Hopital's rule... step by step, and here we go: Refer to my previous message: \[f(x)=\tan \frac{x}{2}\] So your limit is \[\lim_{x \rightarrow 0}\frac{\csc x - \cot x}{x}=\lim_{x \rightarrow 0}\frac{f(x)}{x}\] Getting it so far? The 'tricky' part comes next...

OpenStudy (anonymous):

than?

terenzreignz (terenzreignz):

What is f(0)?

OpenStudy (anonymous):

you mean answer is zero 0?

terenzreignz (terenzreignz):

No, but f(0) is 0, right?

OpenStudy (anonymous):

No. terenzreignz is asking what\[\tan(\frac{0}{2})\]is

OpenStudy (anonymous):

i think its zero?? am i right?

OpenStudy (anonymous):

This is my most hated question and subject

terenzreignz (terenzreignz):

Indeed \[f(0)=\tan\frac{0}{2}=\tan 0 = 0\] So... f(x) = f(x) - f(0) RIGHT? I mean, subtracting zero from a function won't affect it, right?

OpenStudy (anonymous):

yeh

terenzreignz (terenzreignz):

Okay, on another note x = x - 0 This is obvious right?

OpenStudy (anonymous):

Prove: sinx=1+cosx

OpenStudy (anonymous):

yeh its rite but answer is not zero//

terenzreignz (terenzreignz):

sin x is not equal to 1 + cos x. Anyway We have established that f(x) = f(x) - f(0) and x = x - 0 THEREFORE \[\lim_{x \rightarrow 0}\frac{f(x)}{x}=\lim_{x \rightarrow 0}\frac{f(x)-f(0)}{x-0}\] This must look familiar to you by now...

OpenStudy (anonymous):

@terenzreignz it would be nice if u made one clear post. i dont get this problem and i very much would like to

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok than whats the answer?

OpenStudy (anonymous):

tell me the answer please i cant understand it?

terenzreignz (terenzreignz):

Have you taken up derivatives?

OpenStudy (anonymous):

you can just post the worked out solutions you know. later we can ask questions if we dont get any parts...

OpenStudy (raden):

my ideas is : lim (x->0) (csc(x) - cot(x))/x = lim (x->0) [1/sin(x) - cos(x)/sin(x)]/x = lim (x->0) (1- cos(x))/(x*sin(x)) now, u can use the identity : 1-cos(x) = 2 sin^2 (x/2), so it can be = lim (x->0) 2 sin^2 (x/2)/(x*sin(x)) = lim (x->0) 2 sin(x/2) sin(x/2) /(x*sin(x)) and finally, u can use these formulas to find the value that limit lim (x->0) sinAx/Bx = A/B and also lim (x->0) sinAx/sinBx = A/B good luck

OpenStudy (anonymous):

the answer in my book is 1/2 but how?

OpenStudy (raden):

yes, correct. the answer is 1/2

OpenStudy (anonymous):

can you show me the steps please??

OpenStudy (raden):

i have showed u above, u just understanding my steps and i hope u can continue it and to get the answer

terenzreignz (terenzreignz):

This is a summary. We started with \[\lim_{x \rightarrow 0}\frac{\csc x - \cot x}{x}\] Now it can be shown that \[\csc x - \cot x = \tan \frac{x}{2}\] So it becomes \[\lim_{x \rightarrow 0}\frac{\tan\frac{1}{2}x}{x}\] But \[\tan \frac{1}{2}(0) = 0\] meaning it can be subtracted from the numerator without changing its value. Also, 0 can be subtracted from the denominator, without changing its value. The limit then becomes\[\lim_{x \rightarrow 0}\frac{\tan\frac{1}{2}x -\tan\frac{1}{2}(0)}{x - 0}\] But this is just the derivative of tan(x/2) evaluated at x = 0 \[\huge =\frac{d}{dx}\left[ \tan\frac{1}{2}x \right]_{x=0}\] \[\huge \frac{d}{dx}\tan\frac{1}{2}x=\frac{1}{2}\sec^{2}\frac{1}{2}x\] Now just evaluate it at x = 0

OpenStudy (anonymous):

i cant///

OpenStudy (raden):

look the last equation is lim (x->0) 2 sin(x/2) sin(x/2) /(x*sin(x)) = lim (x->0) 2 sin(x/2)/x * sin(x/2)/sinx = 2 * (1/2)/1 * (1/2)/1 = 1/2

OpenStudy (anonymous):

how i cant solve it..

OpenStudy (anonymous):

radEn can you do it more easy to understand>?

OpenStudy (anonymous):

hellow any body here to solve it?

terenzreignz (terenzreignz):

Read the summary.

OpenStudy (anonymous):

i cant

OpenStudy (anonymous):

i cant understand any thing//

OpenStudy (anonymous):

how tan=sec

OpenStudy (anonymous):

\[\tan(x) \neq \sec(x)\]However,\[\frac{d}{dx}\tan(x)=\sec^{2}(x)\]When combined with the chain rule:\[\frac{d}{dx}f(g(x))=f'(g(x)) \times g'(x)\]We can say that\[\frac{d}{dx}\tan(\frac{x}{2})=\frac{1}{2}\sec^{2}(\frac{x}{2})\]

OpenStudy (anonymous):

hahaha i cant understand it again..